In testing the hypothesis that the population mean of a part is 60 inches, a sample of 100 parts resulted in a sample mean of 61.2 inches. The population standard deviation is 7.8 inches. Using a 0.05 level of significance, does this indicate that the population mean is greater than 60 inches?
Answer by ewatrrr(24785)

Ho: u = 60 (claim)
Ha: u ≠ 60
....
sample mean: 61.2
Population: mean = 60in, SD = 7.8in,
Sample n = 100
........
z( 61.2in) = 1.2/(7.8/sqrt(100)) = 1.2/.78 = 1.5385
p-value = P(z > 1.5385) = .06
Using a 0.05 level of significance
.06 > .05
Since the p-value is greater than 5%, fail to reject Ho.
Conclusion: The claim is reasonable at the 5% confidence level.